# Without Paired Labeled Data: End-to-End Self-Supervised Learning for Drone-view Geo-Localization

### v3

@@ III-D Pseudo-Label Enhancement
In self-supervised learning, the quality of pseudo-labels exerts a decisive influence on model performance. Inspired by [49], we propose a pseudo-label optimization strategy based on dynamic feature robustness. Specifically, we introduce small perturbations to the original feature representations and exploit the top-K neighbor consistency between the original and perturbed features to identify high-confidence samples while filtering out noisy instances. Then, a spatial smoothing operation based on the similarity matrix is applied to refine the pseudo-labels to enhance local label consistency.
Small Gaussian noise $\boldsymbol{\epsilon}_{d},\boldsymbol{\epsilon}_{s}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ is used in the initial satellite instances feature and the initial drone instances features to obtain perturbed features, where the $\sigma$ denotes the standard deviation. It can be formulated as
$\mathbf{f}_{i}^{d{\boldsymbol{\epsilon}}}=\mathbf{f}_{i}^{d}+\boldsymbol{\epsilon}_{d},\quad\mathbf{f}_{j}^{s{\boldsymbol{\epsilon}}}=\mathbf{f}_{j}^{s}+\boldsymbol{\epsilon}_{s}$ (34)
@@
$I_{\mathbf{f}_{i}^{d}}=\{y_{\mathbf{f}_{i}^{d}}[k]\cap y_{\mathbf{f}_{i}^{d{\boldsymbol{\epsilon}}}}[k]\}_{k=1}^{N}$ (35)
where $I_{\mathbf{f}_{i}^{d}}$ records the samples with the same identity in instances. Accordingly, we update the cross-view pseudo-label for $\mathbf{f}_{j}^{s}$ using the label with the highest frequency in $I_{\mathbf{f}_{i}^{d}}$, which reflects the most stable identity under both original and perturbed feature rankings. This final refined label is denoted as $\tilde{y}_{\mathbf{f}_{j}^{s}}$. Then, we convert the label list $\{\tilde{y}_{\mathbf{f}_{j}^{s}}\}$ into the form of one-hot code matrix $\tilde{\mathbf{y}}\in\mathbb{R}^{M\times N}$ by setting the column according to the refined labels to 1 and the rest to 0.
In order to explore the intra-view initial and perturbed collaboration, the initial and perturbed similarity matrices $\mathbf{P}^{ss}$ and $\mathbf{P}^{ss}_{\boldsymbol{\epsilon}}$ are calculated, and their sum $\mathbf{P}$ is used to investigate intra-view ranking consistency. Then we keep the 5-max values of $\mathbf{P}$ in each row at 1 and set the rest to 0, acquiring the ranking relations. The process of intra-view ranking smoothing is formulated as
$\mathbf{y}=\mathbf{P}\tilde{\mathbf{y}}$ (36)
where $\mathbf{y}\in\mathbb{R}^{M\times N}$ is the final refined cross-view label matrix of satellite instance. In $\mathbf{y}$, the column number of the maximum value in each row is the refined label of the samples.
TABLE I: Comparisons between the proposed method and some state-of-the-art methods on the University-1652 datasets.
@@
Model Venue Supervision Ratio Parameters(M) R@1 AP R@1 AP
MuSe-Net[52] PR’2024 100(%) 50.47 74.48 77.83 88.02 75.10
MCCG[1] TCSVT’2023 100(%) 56.65 89.40 91.07 95.01 89.93
SDPL[53] TCSVT’2024 100(%) 42.56 90.16 91.64 93.58 89.45
CCR[31] TCSVT’2024 100(%) 156.57 92.54 93.78 95.15 91.80
Sample4Geo[28] ICCV’2023 100(%) 87.57 92.65 93.81 95.14 91.39
SRLN[30] TGRS’2024 100(%) 193.03 92.70 93.77 95.14 91.97
MEAN[2] arXiv’2024 100(%) 36.50 93.55 94.53 96.01 92.08
DAC[3] TCSVT’2024 100(%) 96.50 94.67 95.50 96.43 93.79
Ours - 100(%) 28.59 92.02 93.36 95.57 91.27
EM-CVGL[4] TGRS’2024 0(%) 2.25 70.29 74.93 79.03 61.03
CDIKTNet[40] arXiv’2025 0(%) 91.52 83.30 85.73 87.73 76.53
Wang[9] AAAI’2025 0(%) - 85.95 90.33 94.01 82.66
Ours - 0(%) 28.59 90.17 91.67 95.01 88.95
@@ IV-A Experimental Datasets and Evaluation Metrics
To evaluate the proposed method, we conduct experiments on three benchmark datasets, including University-1652 [12], SUES-200 [32] and DenseUAV [52].
University-1652 contains abundant images of 1,652 buildings in 72 universities around the world. It involves drone, satellite, and ground-level views. The training set consists of 701 buildings from 33 universities and the test set covers 951 buildings from 39 geographically nonaffiliated universities.
SUES-200 focuses on altitude diversity in aerial imagery, containing 200 locations, with 120 designated for training and 80 for testing. Each location in the dataset contributes one satellite image along with aerial images captured at four distinct altitudes, which involve urban areas, natural landscapes, and water bodies. SUES-200 evaluates cross-view retrieval under varying flight heights, reflecting real-world drone scenarios.
DenseUAV focuses on drone self-positioning in low-altitude urban environments with more than 27,000 drone and satellite images collected from 14 university campuses. It is characterized by dense sampling, multi-scale satellite imagery, and diverse altitudes captured across multiple temporal conditions. DenseUAV poses challenges such as cross-view matching, neighbor confusion, and spatial-temporal variation.
We evaluate the models using Recall@K (R@K) and Average Precision (AP), where R@K measures the proportion of correct matches among the Top-$K$results and AP evaluates the overall precision-recall performance. Model efficiency is assessed by the number of parameters to reflect portability under resource constraints. All comparisons are conducted using the best configurations for each method.
IV-B Implementation Details
A ConvNeXt-Tiny model, pre-trained on ImageNet, serves as the backbone for feature extraction, with a newly added classifier module initialized via the Kaiming method. During both training and testing, all input images are resized to $3\times 384\times 384$. We employ several data-augmentation techniques, including random cropping, random horizontal flipping, and random rotation. The batch size is set to 64, and DBSCAN [7] is used to generate pseudo-labels. For optimization, we use the SGD optimizer with an initial learning rate of 0.001, training the model for 30 epochs in total. All experiments are conducted under the PyTorch framework on an Ubuntu 22.04 system equipped with four NVIDIA RTX 4090 GPUs. Additionally, for the University-1652 and SUES-200 datasets, we apply feature replication only on the training sets by duplicating each satellite feature 50 times for clustering purposes, while drone features remain unchanged. The test sets are kept intact without any modification. In the DenseUAV dataset, drone features are replicated four times at each altitude, and satellite features are replicated three times for each time point and altitude exclusively in the training set; the test set remains unchanged.
IV-C Comparison with State-of-the-art Methods
To evaluate the effectiveness of our method, we compare it with the state-of-the-art supervised methods MuSe-Net [52], MCCG [1], SDPL [53], CCR [31], Sample4Geo [28], SRLN [30], MEAN [2] and DAC [3], and the self-supervised methods EM-CVGL [4], CDIKTNet [40] and Wang [9].
[…]
We further benchmark with eight representative supervised methods, and the results show that our method clearly surpasses some of them, including MuSe-Net, MCCG, and SDPL. Furthermore, with self-supervised pre-training and subsequent supervised fine-tuning, our method achieves competitive performance under the supervised setting, comparable to CCR, Sample4Geo, and SRLN.
[…]
@@ IV-C Comparison with State-of-the-art Methods
150m 200m 250m 300m
Model Venue Supervision Ratio Parameters(M) R@1 AP R@1 AP R@1 AP R@1 AP
MCCG[1] TCSVT’2023 100(%) 56.65 82.22 85.47 89.38 91.41 93.82 95.04 95.07 96.20
SDPL[53] TCSVT’2024 100(%) 42.56 82.95 85.82 92.73 94.07 96.05 96.69 97.83 98.05
CCR[31] TCSVT’2024 100(%) 156.57 87.08 89.55 93.57 94.90 95.42 96.28 96.82 97.39
SRLN[30] TGRS’2024 100(%) 193.03 89.90 91.90 94.32 95.65 95.92 96.79 96.37 97.21
Sample4Geo[28] ICCV’2023 100(%) 87.57 92.60 94.00 97.38 97.81 98.28 98.64 99.18 99.36
DAC[3] TCSVT’2024 100(%) 96.50 96.80 97.54 97.48 97.97 98.20 98.62 97.58 98.14
MEAN[2] arXiv’2024 100(%) 36.50 95.50 96.46 98.38 98.72 98.95 99.17 99.52 99.63
Ours - 100(%) 28.59 94.00 95.21 97.83 98.88 99.14 95.44 99.25 99.39
EM-CVGL[4] TGRS’2024 0(%) 2.25 55.23 60.80 60.95 61.03 68.10 72.62 74.42 78.20
CDIKTNet[40] arXiv’2025 0(%) 91.52 82.75 85.25 89.35 91.13 93.15 94.39 95.18 96.12
Wang[9] AAAI’2025 0(%) - 76.90 84.95 87.88 92.60 92.98 95.66 95.10 96.92
Ours - 0(%) 28.59 82.60 85.60 90.48 92.28 94.33 95.44 96.78 97.42
Ours - 0(%) 28.59 87.00 89.20 93.98 95.11 96.50 97.23 97.23 97.86
Ours - 0(%) 28.59 90.98 92.66 95.03 96.14 97.45 98.03 98.23 98.64
@@
150m 200m 250m 300m
Model Venue Supervision Ratio Parameters(M) R@1 AP R@1 AP R@1 AP R@1 AP
CCR[31] TCSVT’2024 100(%) 156.57 92.50 88.54 97.50 95.22 97.50 97.10 97.50 97.49
Sample4Geo[28] ICCV’2023 100(%) 87.57 92.60 94.00 97.38 97.81 98.28 98.64 99.18 99.36
MCCG[1] TCSVT’2023 100(%) 56.65 93.75 89.72 93.75 92.21 96.25 96.14 98.75 96.64
SDPL[53] TCSVT’2024 100(%) 42.56 93.75 83.75 96.25 92.42 97.50 95.65 96.25 96.17
SRLN[30] TGRS’2024 100(%) 193.03 93.75 93.01 97.50 95.08 97.50 96.52 97.50 96.71
DAC[3] TCSVT’2024 100(%) 96.50 97.50 94.06 98.75 96.66 98.75 98.09 98.75 97.87
MEAN[2] arXiv’2024 100(%) 36.50 97.50 94.75 100.00 97.09 100.00 98.28 100.00 99.21
Ours - 100(%) 28.59 97.50 92.71 97.50 97.43 100.00 98.98 98.75 98.25
EM-CVGL[4] TGRS’2024 0(%) 2.25 73.75 54.00 91.25 65.65 96.25 72.02 97.50 74.74
CDIKTNet[40] arXiv’2025 0(%) 91.52 88.75 80.33 93.75 88.17 95.00 92.15 98.75 94.37
Wang[9] AAAI’2025 0(%) - 87.50 74.81 92.50 87.15 96.25 91.20 98.75 94.52
Ours - 0(%) 28.59 91.25 81.28 97.50 90.62 98.75 94.74 97.50 96.17
Ours - 0(%) 28.59 95.00 85.46 97.50 92.72 100.00 95.43 100.00 96.39
Ours - 0(%) 28.59 96.25 90.05 97.50 94.80 98.75 96.82 98.75 97.33
Results on SUES-200. As shown in Table II, our method, when directly transferring the pre-trained model from the University-1652 dataset to the SUES-200 dataset without any re-training, consistently outperforms all existing self-supervised methods and achieves competitive performance compared to supervised methods such as MCCG and SDPL. Moreover, by performing self-supervised re-training on the SUES-200 dataset, our method brings a significant performance boost, achieving results comparable to the supervised CCR. Notably, when further fine-tuning the transferred model for only one epoch, we observe additional improvements, surpassing the supervised SRLN. These results clearly demonstrate the effectiveness of our method in self-supervised learning, transfer learning, and efficient fine-tuning.
Results on Cross-Domain Generalization Performance. As shown in Table III, our method consistently outperforms existing self-supervised methods in cross-domain settings, showing both significant improvements and stable generalization. Specifically, in the Drone to Satellite setting, our model achieves better performance than EM-CVGL at all altitudes. At 150 meters, R@1 and AP increase by approximately 22% and 20%. At 200 meters, the gains are around 19% and 16%; at 250 meters, nearly 14% and 12%. Even at 300 meters, our method still outperforms by about 11% and 9%. In the Satellite to Drone scenario, similar improvements are observed. At 150 meters, R@1 and AP improve by approximately 18% and 25%; at 200 meters, by 10% and 20%; at 250 meters, by 6% and 13%; and at 300 meters, by 3% and 10%. These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes. Remarkably, our method even surpasses all existing state-of-the-art supervised methods in terms of cross-domain generalization. These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes.
TABLE III: Comparisons between the proposed method and state-of-the-art methods in cross-domain evaluation.
@@
150m 200m 250m 300m
Model Venue Supervision Ratio Parameters(M) R@1 AP R@1 AP R@1 AP R@1 AP
MCCG[1] TCSVT’2023 100(%) 56.65 57.62 62.80 66.83 71.60 74.25 78.35 82.55 85.27
Sample4Geo[28] ICCV’2023 100(%) 87.57 70.05 74.93 80.68 83.90 87.35 89.72 90.03 91.91
DAC[3] TCSVT’2024 100(%) 96.50 76.65 80.56 86.45 89.00 92.95 94.18 94.53 95.45
CAMP[29] TGRS’2024 100(%) 91.40 78.90 82.38 86.83 89.28 91.95 93.63 95.68 96.65
MEAN[2] arXiv’2024 100(%) 36.50 81.73 87.72 89.05 91.00 92.13 93.60 94.63 95.76
EM-CVGL[4] TGRS’2024 0(%) 2.25 60.03 65.69 71.50 76.17 80.35 83.85 85.93 88.34
Ours - 0(%) 28.59 82.60 85.60 90.48 92.28 94.33 95.44 96.78 97.42
[…]
150m 200m 250m 300m
Model Venue Supervision Ratio Parameters(M) R@1 AP R@1 AP R@1 AP R@1 AP
MCCG[1] TCSVT’2023 100(%) 56.65 61.25 53.51 82.50 67.06 81.25 74.99 87.50 80.20
Sample4Geo[28] ICCV’2023 100(%) 87.57 83.75 73.83 91.25 83.42 93.75 89.07 93.75 90.66
DAC[3] TCSVT’2024 100(%) 96.50 87.50 79.87 96.25 88.98 95.00 92.81 96.25 94.00
CAMP[29] TGRS’2024 100(%) 91.40 87.50 78.98 95.00 87.05 95.00 91.05 96.25 93.44
MEAN[2] arXiv’2024 100(%) 36.50 91.25 81.50 96.25 89.55 95.00 92.36 96.25 94.32
EM-CVGL[4] TGRS’2024 0(%) 2.25 73.75 56.99 87.50 70.62 92.50 81.18 95.00 86.04
Ours - 0(%) 28.59 91.25 81.28 97.50 90.62 98.75 94.74 97.50 96.17
TABLE IV: Comparisons between the proposed method and some state-of-the-art methods on the DenseUAV datasets.
DenseUAV
All height 80m 90m 100m
Model Venue Supervision Ratio Parameters(M) R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5
ConvNext-Tiny[33] TIP’2024 100(%) 30.10 60.23 81.94 - - - - - -
Sample4Geo[28] ICCV’2023 100(%) 87.57 80.57 96.53 80.57 96.53 86.87 98.71 91.38 99.74
CAMP[29] TGRS’2024 100(%) 91.40 88.72 97.76 76.45 96.01 84.81 99.23 89.19 99.23
DAC[3] TCSVT’2024 100(%) 96.50 84.47 96.53 80.92 97.04 85.32 98.71 85.46 99.10
MEAN[2] arXiv’2024 100(%) 36.50 90.18 97.86 78.76 97.04 85.33 99.10 89.32 98.97
Ours - 100(%) 28.59 86.31 97.81 86.10 98.46 88.93 99.61 90.35 99.49
EM-CVGL[4] TGRS’2024 0(%) 2.25 18.15 50.97 18.02 50.36 17.76 50.19 17.50 49.98
CDIKTNet[40] arXiv’2025 0(%) 91.52 22.57 40.24 17.25 40.15 23.04 48.39 25.48 52.38
Ours - 0(%) 28.59 73.79 90.69 73.10 92.92 74.65 95.37 75.03 95.50
Results on DenseUAV. The DenseUAV dataset presents a highly challenging scenario due to its densely sampled nature, where neighboring images exhibit high visual similarity, often leading to severe confusion in DVGL. In addition, the dataset includes satellite images captured at different times and drone images from multiple altitudes, which further increases intra-domain variability and task complexity. As shown in Table IV, in the fully unsupervised setting, our method achieves 73.79% R@1 and 90.69% R@5 across all altitudes, significantly outperforming existing self-supervised methods. It is noteworthy that the two representative self-supervised baselines consistently yield poor performance on DenseUAV, due to their limitations in handling densely sampled and spatio-temporal diverse scenarios. Furthermore, after supervised fine-tuning, our method surpasses several supervised baselines, including ConvNeXt-Tiny, Sample4Geo, CAMP, and DAC. In some cases, it even outperforms the current best-performing supervised method, MEAN. These results clearly demonstrate the effectiveness of our method in addressing extreme feature confusion caused by spatial continuity and appearance shifts induced by temporal variation, confirming its strong generalization and robustness.

### v5

@@ III-D Pseudo-Label Enhancement
In self-supervised learning, the quality of pseudo-labels exerts a decisive influence on model performance. Inspired by [yang2024shallow], we propose a pseudo-label optimization strategy based on dynamic feature robustness. Specifically, we introduce small perturbations to the original feature representations and exploit the top-K neighbor consistency between the original and perturbed features to identify high-confidence samples while filtering out noisy instances. Then, a spatial smoothing operation based on the similarity matrix is applied to refine the pseudo-labels to enhance local label consistency.
Small Gaussian noise $\boldsymbol{\epsilon}_{d},\boldsymbol{\epsilon}_{s}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ is used in the initial satellite instances feature and the initial drone instances features to obtain perturbed features, where $\sigma$ denotes the standard deviation. It can be formulated as
$\mathbf{f}_{i}^{d{\boldsymbol{\epsilon}}}=\mathbf{f}_{i}^{d}+\boldsymbol{\epsilon}_{d},\quad\mathbf{f}_{j}^{s{\boldsymbol{\epsilon}}}=\mathbf{f}_{j}^{s}+\boldsymbol{\epsilon}_{s}.$ (24)
@@
$I_{\mathbf{f}_{i}^{d}}=\{y_{\mathbf{f}_{i}^{d}}[k]\cap y_{\mathbf{f}_{i}^{d{\boldsymbol{\epsilon}}}}[k]\}_{k=1}^{N},$ (25)
where $I_{\mathbf{f}_{i}^{d}}$ records the samples with the same identity in instances. Consequently, we update the cross-view pseudo-label for $\mathbf{f}_{j}^{s}$ using the label with the highest frequency in $I_{\mathbf{f}_{i}^{d}}$, which reflects the most stable identity under both original and perturbed feature rankings. This final refined label is denoted as $\tilde{y}_{\mathbf{f}_{j}^{s}}$. Then, we convert the label list $\{\tilde{y}_{\mathbf{f}_{j}^{s}}\}$ into the form of one-hot code matrix $\tilde{\mathbf{y}}\in\mathbb{R}^{M\times N}$ by setting the column according to the refined labels to 1 and the rest to 0.
In order to explore the intra-view initial and perturbed collaboration, the initial and perturbed similarity matrices $\mathbf{P}^{ss}$ and $\mathbf{P}^{ss}_{\boldsymbol{\epsilon}}$ are calculated, and their sum $\mathbf{P}$ is used to investigate intra-view ranking consistency. Then we keep the 5-max values of $\mathbf{P}$ in each row at 1 and set the rest to 0, acquiring the ranking relations. The process of intra-view ranking smoothing is formulated as $\mathbf{y}=\mathbf{P}\tilde{\mathbf{y}}$, where $\mathbf{y}\in\mathbb{R}^{M\times N}$ is the final refined cross-view label matrix of the satellite instance. In $\mathbf{y}$, the column number of the maximum value in each row is the refined label of the samples.
TABLE I: Comparisons between DMNIL and existing methods on University-1652
@@
Model Learning Type Parameters(M) R@1 AP R@1 AP
MuSe-Net[wang2024multiple] Supervised 50.47 74.48 77.83 88.02 75.10
MCCG[shen2023mccg] Supervised 56.65 89.40 91.07 95.01 89.93
SDPL[chen2024sdpl] Supervised 42.56 90.16 91.64 93.58 89.45
Sample4Geo[deuser2023sample4geo] Supervised 87.57 92.65 93.81 95.14 91.39
SRLN[lv2024direction] Supervised 193.03 92.70 93.77 95.14 91.97
MEAN[chen2024multi] Supervised 36.50 93.55 94.53 96.01 92.08
DAC[xia2024enhancing] Supervised 96.50 94.67 95.50 96.43 93.79
EM-CVGL[li2024learning] Self-Supervised 2.25 70.29 74.93 79.03 61.03
CDIKTNet[chen2025limited] Self-Supervised 91.52 83.30 85.73 87.73 76.53
Wang[wang2025coarse] Self-Supervised - 85.95 90.33 94.01 82.66
DMNIL Self-Supervised 28.59 90.17 91.67 95.01 88.95
@@ IV-A Experimental Datasets and Evaluation Metrics
To evaluate the proposed method, we conduct experiments on three benchmark datasets, including University-1652 [zheng2020university], SUES-200 [zhu2023sues] and DenseUAV [wang2024multiple].
University-1652 contains multi-view images of 1,652 buildings across 72 universities worldwide, including drone, satellite, and ground-level views. The training set includes 701 buildings from 33 universities, and the test set covers 951 buildings from 39 geographically distinct universities. SUES-200 emphasizes altitude diversity, consisting of 200 locations (120 for training and 80 for testing). Each location provides one satellite image and four aerial images captured at different altitudes, covering urban, natural, and aquatic scenes. DenseUAV focuses on low-altitude drone self-positioning, with over 27,000 drone–satellite images from 14 university campuses. It features dense sampling, multi-scale satellite imagery, and diverse temporal conditions, presenting challenges such as cross-view misalignment and spatial–temporal variations.
We evaluate the models using Recall@K (R@K) and Average Precision (AP), where R@K measures the proportion of correct matches among the Top-KKresults and AP evaluates the overall precision-recall performance. Model efficiency is assessed by the number of parameters to reflect portability under resource constraints. All comparisons are conducted using the best configurations for each method.
TABLE II: Comparisons between DMNIL and existing methods on SUES-200. $\dagger$} indicates results from self-supervised training on University-1652 and testing on SUES-200. $\ddagger$} denotes results from self-supervised training and evaluation on SUES-200. $\ast$} corresponds to results after one epoch self-supervised fine-tuning on SUES-200 based on the $\dagger$} pre-trained model
@@
150m 200m 250m 300m 150m 200m 250m 300m
Model Learning Type Parameters(M) R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP
MCCG[shen2023mccg] Supervised 56.65 82.22 85.47 89.38 91.41 93.82 95.04 95.07 96.20 93.75 89.72 93.75 92.21 96.25 96.14 98.75 96.64
SDPL[chen2024sdpl] Supervised 42.56 82.95 85.82 92.73 94.07 96.05 96.69 97.83 98.05 93.75 83.75 96.25 92.42 97.50 95.65 96.25 96.17
SRLN[lv2024direction] Supervised 193.03 89.90 91.90 94.32 95.65 95.92 96.79 96.37 97.21 93.75 93.01 97.50 95.08 97.50 96.52 97.50 96.71
Sample4Geo[deuser2023sample4geo] Supervised 87.57 92.60 94.00 97.38 97.81 98.28 98.64 99.18 99.36 97.50 93.63 98.75 96.70 98.75 98.28 98.75 98.05
DAC[xia2024enhancing] Supervised 96.50 96.80 97.54 97.48 97.97 98.20 98.62 97.58 98.14 97.50 94.06 98.75 96.66 98.75 98.09 98.75 97.87
MEAN[chen2024multi] Supervised 36.50 95.50 96.46 98.38 98.72 98.95 99.17 99.52 99.63 97.50 94.75 100.00 97.09 100.00 98.28 100.00 99.21
DMNIL Supervised 28.59 94.00 95.21 97.83 98.88 99.14 95.44 99.25 99.39 97.50 92.71 97.50 97.43 100.00 98.98 98.75 98.25
EM-CVGL[li2024learning] Self-Supervised 2.25 55.23 60.80 60.95 61.03 68.10 72.62 74.42 78.20 73.75 54.00 91.25 65.65 96.25 72.02 97.50 74.74
CDIKTNet[chen2025limited] Self-Supervised 91.52 82.75 85.25 89.35 91.13 93.15 94.39 95.18 96.12 88.75 80.33 93.75 88.17 95.00 92.15 98.75 94.37
Wang[wang2025coarse] Self-Supervised - 76.90 84.95 87.88 92.60 92.98 95.66 95.10 96.92 87.50 74.81 92.50 87.15 96.25 91.20 98.75 94.52
DMNIL Self-Supervised 28.59 82.60 85.60 90.48 92.28 94.33 95.44 96.78 97.42 91.25 81.28 97.50 90.62 98.75 94.74 97.50 96.17
DMNIL Self-Supervised 28.59 87.00 89.20 93.98 95.11 96.50 97.23 97.23 97.86 95.00 85.46 97.50 92.72 100.00 95.43 100.00 96.39
DMNIL Self-Supervised 28.59 90.98 92.66 95.03 96.14 97.45 98.03 98.23 98.64 96.25 90.05 97.50 94.80 98.75 96.82 98.75 97.33
IV-B Implementation Details
A ConvNeXt-Tiny model, pre-trained on ImageNet, serves as the backbone for feature extraction, with a newly added classifier module initialized via the Kaiming method. During both training and testing, all input images are resized to $3\times 384\times 384$. We employ several data-augmentation techniques, including random cropping, random horizontal flipping, and random rotation. The batch size is set to 64. We use the DBSCAN [ester1996density] to generate pseudo-labels, where the neighborhood radius is set to 0.40 for drone-view features and 0.30 for satellite-view features, and the minimum number of points required to form a cluster is 4. For optimization purposes, we use the SGD optimizer with an initial learning rate of 0.001 to train the model for 30 epochs in total. The weight $w{{}^{\prime\prime}}$ and $w{{}^{\prime}}$ in Eq.(11) are set to 0.7 and 0.3, respectively. The $\gamma$ in Eq.(16) is set to 0.9. The top-$k_{1}$ and top-$k_{2}$ are set to 10 and 20, respectively. The $\lambda_{0}$, $\lambda_{1}$, and $\lambda_{2}$ are set to 0.2, 0.01, and 0.1, respectively. All experiments are conducted under the PyTorch framework on an Ubuntu 22.04 system equipped with four NVIDIA RTX 4090 GPUs.
IV-C Comparison with State-of-the-art Methods
To evaluate the effectiveness of DMNIL, we compare it with the state-of-the-art supervised methods MuSe-Net [wang2024multiple], MCCG[shen2023mccg], SDPL [chen2024sdpl], Sample4Geo [deuser2023sample4geo], SRLN [lv2024direction], MEAN [chen2024multi] and DAC [xia2024enhancing], and the self-supervised methods EM-CVGL [li2024learning], CDIKTNet [chen2025limited] and Wang [wang2025coarse]. Notably, self-supervised solutions for DVGL remain limited, and the development of effective self-supervised methods in this domain is still in its early stage. This highlights our contributions to the advancement of self-supervised learning for DVGL.
@@ IV-C Comparison with State-of-the-art Methods
Results on SUES-200. As shown in Table II, when directly transferring the pre-trained model from the University-1652 dataset to the SUES-200 dataset without any re-training, DMNIL consistently outperforms all existing self-supervised methods and achieves competitive performance compared with supervised methods such as SDPL. DMNIL outperforms some supervised methods such as MCCG and SDPL. Notably, when further fine-tuning the transferred model for only one epoch, we observe additional improvements that surpass the supervised SRLN. These results clearly demonstrate the effectiveness of DMNIL in self-supervised learning, transfer learning, and efficient fine-tuning.
Results on Cross-Domain Generalization Performance. As shown in Table III, DMNIL consistently outperforms existing self-supervised methods in cross-domain setting. Specifically, in the Drone $\rightarrow$ Satellite setting, our model achieves better performance than EM-CVGL at all altitudes. At 150 meters, R@1 and AP increase by approximately 22% and 20%. At 200 meters, the lifts from DMNIL are around 19% and 16%. At 250 meters, they are nearly 14% and 12%. Even at 300 meters, DMNIL still outperforms about 11% and 9%. In the Satellite $\rightarrow$ Drone scenario, similar improvements are observed. At 150 meters, R@1 and AP increase by approximately 18% and 25%, respectively. At 200 meters, the improvements are around 10% and 20%. Moving further to 250 meters, they are lifted by 6% and 13%. Finally, at 300 meters, the lifts remain considerable with R@1 increasing by 3% and AP increasing by 10%. Remarkably, DMNIL even surpasses all existing state-of-the-art supervised methods in terms of cross-domain generalization. These results confirm the robustness and strong cross-domain generalization capability of DMNIL at varying altitudes.
TABLE III: Comparisons between DMNIL and existing methods in cross-domain evaluation
@@
150m 200m 250m 300m 150m 200m 250m 300m
Model Learning Type Parameters(M) R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP
MCCG[shen2023mccg] Supervised 56.65 57.62 62.80 66.83 71.60 74.25 78.35 82.55 85.27 61.25 53.51 82.50 67.06 81.25 74.99 87.50 80.20
Sample4Geo[deuser2023sample4geo] Supervised 87.57 70.05 74.93 80.68 83.90 87.35 89.72 90.03 91.91 83.75 73.83 91.25 83.42 93.75 89.07 93.75 90.66
CAMP[wu2024camp] Supervised 91.40 78.90 82.38 86.83 89.28 91.95 93.63 95.68 96.65 87.50 78.98 95.00 87.05 95.00 91.05 96.25 93.44
DAC[xia2024enhancing] Supervised 96.50 76.65 80.56 86.45 89.00 92.95 94.18 94.53 95.45 87.50 79.87 96.25 88.98 95.00 92.81 96.25 94.00
MEAN[chen2024multi] Supervised 36.50 81.73 87.72 89.05 91.00 92.13 93.60 94.63 95.76 91.25 81.50 96.25 89.55 95.00 92.36 96.25 94.32
EM-CVGL[li2024learning] Self-Supervised 2.25 60.03 65.69 71.50 76.17 80.35 83.85 85.93 88.34 73.75 56.99 87.50 70.62 92.50 81.18 95.00 86.04
DMNIL Self-Supervised 28.59 82.60 85.60 90.48 92.28 94.33 95.44 96.78 97.42 91.25 81.28 97.50 90.62 98.75 94.74 97.50 96.17
Results on DenseUAV. The DenseUAV dataset presents a highly challenging scenario due to its densely sampled nature, where neighboring images exhibit high visual similarity. It often leads to severe confusion in DVGL. In addition, the dataset includes satellite images captured at different times and drone images from multiple altitudes, which further increases intra-domain variability and task complexity. As shown in Table IV, in the fully self-supervised setting, DMNIL achieves 73.79% R@1 and 90.69% R@5 across all altitudes, outperforming existing self-supervised methods. It is noteworthy that the two representative self-supervised methods consistently yield poor performance on DenseUAV, due to their limitations in handling densely sampled and spatio-temporal diverse scenarios. These results demonstrate the effectiveness of DMNIL in addressing feature confusion caused by spatial continuity and appearance shifts induced by temporal variation.